# Analysis methodology of neutron time-of-flight spectra based on Bayesian unfolding and accurate Monte Carlo simulations

### v1

@@ 1 Introduction
  • a high intrinsic neutron detection efficiency ($\sim 30$ %);
@@
$TOF=R\cdot E,$ (2)
where here $TOF$ is the TOF spectrum, $E$ is the neutron energy distribution, and $R$ is the response matrix of the detection system.
@@ 1.2 TOF spectrum unfolding
  • $\sum_{i=1}^{n_{C}}P(C_{i}|E_{j})=1$—that is, any observed effect must come from one or more of the considered causes;
@@ 2 Monte Carlo simulations
As mentioned in Section 1.2, the response matrix $P(E_{j}|C_{i})$ can be calculated through Monte Carlo simulations. In this work, the Geant4 simulation toolkit [21] was used to determine the TOF response of the entire experimental setup to individual neutron energy ranges. Due to the complex nature of neutron interactions, a good physics model and correct implementation of the detector geometries and surrounding materials are very important for calculating accurate TOF distributions. For this reason, the G4ParticleHP model [22] in Geant4 was used, since it has been validated extensively for the simulation of MONSTER modules [7, 23, 24]. About the geometric model, Figure 2 shows the high level of detail with which MONSTER was implemented in the simulation, with modules placed at $1.5$ m and $2$ m flight paths from the source. For simplicity, the analysis discussed in this paper considers only the MONSTER array at a flight path of $2$ m.
@@
Figure 2: Geometric model of the experimental setup simulated with Geant4. Panel (a) shows the full experimental setup. Panel (b) shows the detailed geometry of a MONSTER module.
Figure 3: MONSTER simulated TOF response to 22 MeV neutrons for several setups, including more or fewer geometry elements in the simulation.
@@ 3.1 Impact of the binning of the response matrix
Figure 5 shows the response matrices with three different types of binning. The vertical axis corresponds to the causes, with a cause being a given range of neutron energies. The horizontal axis corresponds to the effects—that is, the TOF. Several values of bin widths were studied for each type of binning but, for clarity, only the most relevant are presented:
@@
The simulated 85As TOF spectrum was then unfolded with each of these three response matrices. For simplicity, in this first analysis only the neutron emission part of the $\beta$-decay was considered and the effect of the $\beta$-detection efficiency was ignored. The reconstructed TOF spectra obtained with each kind of binning can be seen in Figure 6. They are almost indistinguishable from the simulated ones.
@@ 3.2 Impact of the neutron detection threshold
The effect of the neutron detection threshold, as in the case of a real experiment, has also been studied by analyzing the virtual experiment. The same threshold was applied to the simulated neutron detector pulse height spectrum for producing the TOF spectrum and to the response matrix used in the unfolding. In this process, a realistic non-proportional light output of the liquid scintillators was used in the simulations. The effect of the neutron detection threshold is two-fold: it reduces the number of causes producing detectable effects, and the causes that produce them do it over a smaller TOF range. Figure 8 shows the simulated and reconstructed TOF spectra (top panel) and the resulting neutron energy distribution (bottom panel) after applying a detection threshold of $30$ keVee, corresponding to a neutron energy of $\sim 300$ keV. The value corresponds to the real threshold of MONSTER in the experiment. As can be seen in the figure, the inclusion of the threshold limits the lower neutron energy that can be detected, and hence the original neutron spectrum is never going to be reproduced below that limit. The reconstructed energy distribution has been normalized to the estimated true total number of events $\hat{N}_{true}$, as calculated from Equation 5, and the total number of decays $N_{obs}$.
@@ 4 Conclusions
The methodology discussed in this paper has been applied successfully to the analysis of the $\beta$-delayed neutron energy spectra emitted in the decays of 85,86As. These measurements served as an experimental validation of the methodology and for producing new physics results. These results will be published in a forthcoming paper.

### v2

# Analysis of neutron time-of-flight spectra with a Bayesian unfolding methodology

@@ 1 Introduction
  • a high intrinsic neutron detection efficiency ($\sim 30$);
@@
$\boldsymbol{TOF}=\boldsymbol{R}\cdot\boldsymbol{E},$ (2)
where here $\boldsymbol{TOF}$ is the TOF spectrum, $\boldsymbol{E}$ is the neutron energy distribution, and $\boldsymbol{R}$ is the response matrix of the detection system.
@@ 1.2 TOF spectrum unfolding
  • $\sum_{i=1}^{n_{C}}P(C_{i}|E_{j})=1$, meaning that any observed effect must come from one or more of the considered causes;
@@ 2 Monte Carlo simulations
As mentioned in Section 1.2, the response matrix $P(E_{j}|C_{i})$ can be calculated through Monte Carlo simulations. In this work, the Geant4 simulation toolkit [21] was used to determine the TOF response of the entire experimental setup to individual neutron energy ranges. Due to the complex nature of neutron interactions, a good physics model and correct implementation of the detector geometries and surrounding materials are very important for calculating accurate TOF distributions. For this reason, the G4ParticleHP model [22] in Geant4 was used, since it has been validated extensively for the simulation of MONSTER modules [7, 23, 24]. This high-precision model includes the necessary models and data for a complete description of neutron-induced alpha production reactions on carbon.
The detailed simulations performed in this work include the parametrized light output functions ($L$, in MeVee) of the scintillators for electrons, protons, and heavier ions as a function of their energy ($E$, in MeV). For electrons, the light output function was taken from Reference [25]. For protons and heavier ions, the light output functions were selected from the wide set of functions available in the data files of NRESP7.1 [26], and adapted for the simulation of MONSTER modules by comparing the results of simulations against experimental measurements of response functions for neutrons of different energies:
$\begin{array}[]{ll}L_{p}&=-1.4177+0.68639E_{p}\qquad E_{p}>8\textrm{ MeV}\\ L_{d}&=2L_{p}(0.5E_{d})\\ L_{\alpha}&=\begin{cases}0.02076E_{\alpha}^{1.871}&E_{\alpha}<6.76\textrm{ MeV}\\ -0.70929+0.21454E_{\alpha}&E_{\alpha}\geq 6.76\textrm{ MeV}\end{cases}\\ L_{Be}&=0.01E_{Be}\\ L_{B,C}&=0.0097E_{B,C}.\end{array}$ (7)
For protons below $8$ MeV, the light yield is calculated by linear interpolation of a set of discrete experimental values [27].
In the simulation, all secondary particles giving rise to light in the scintillator are tracked, and the light yield is calculated at the end of each step using the corresponding light output function. The total light yield for each event is obtained by the addition of the light produced in each step by each kind of secondary particle produced. In this way, the simulations were able to reproduce the experimental response functions within a $5$ of uncertainty. Including the light production mechanisms of the scintillators in the simulation allows to take into account important effects such as the neutron detection threshold of the detectors due to the produced light, which in turn allows to calculate the neutron detection efficiency with high accuracy.
About the geometric model, Figure 2 shows the high level of detail with which MONSTER was implemented in the simulation, with modules placed at $1.5$ m and $2$ m flight paths from the source. For simplicity, the analysis discussed in this paper considers only the MONSTER array at a flight path of $2$ m.
@@
Figure 2: Geometric model of the experimental setup simulated with Geant4. Panel (a) shows the full experimental setup, with 30 MONSTER modules at 2 m, 18 MONSTER modules at 1.5 m, and some ancillary detectors surrounding the implantation point. Panel (b) shows the detailed geometry of a MONSTER module: the aluminum cover and supports, and the mu-metal shielding in different shades of gray; the BC501A liquid scintillator in green; in salmon the quartz window; in blue the PMMA light guide; the thin layers of reflective paint in blue and yellow; and in magenta the BC501A expansion volume.
Figure 3: MONSTER simulated TOF response to $2$ MeV neutrons for several setups, including more or fewer geometry elements in the simulation.
@@ 3.1 Impact of the binning of the response matrix
Figure 5 shows the response matrices with three different types of binning. The vertical axis corresponds to the causes. In this case, each cause corresponds to a range of neutron energies. The horizontal axis corresponds to the effects—that is, the TOF. Several values of bin widths were studied for each type of binning but, for clarity, only the most relevant are presented:
@@
The simulated 85As TOF spectrum was then unfolded with each of these three response matrices. For simplicity, in this first analysis only the neutron emission part of the $\beta$-decay was considered and the effect of the $\beta$-detection efficiency was ignored. The reconstructed TOF spectra obtained with each kind of binning can be seen in Figure 6. They are almost indistinguishable from the simulated ones. In addition to the simulated and reconstructed TOF spectra, the contribution of each individual cause is also shown in the figure. This allows to see that the simulated TOF spectrum can be reconstructed accurately with different sets of responses. Furthermore, this shows that peak-like structures in the TOF spectrum do not necessarily correspond to single transitions, but rather to regions of increased level density or feeding.
@@ 3.2 Impact of the neutron detection threshold
(a) Reconstruction
(b) Energy distribution
(a) Reconstruction
(b) Energy distribution
The effect of the neutron detection threshold, as in the case of a real experiment, has also been studied by analyzing the virtual experiment. The same threshold was applied to the simulated neutron detector pulse height spectrum for producing the TOF spectrum and to the response matrix used in the unfolding. In this process, a realistic non-proportional light output of the liquid scintillators was used in the simulations. The effect of the neutron detection threshold is two-fold: it reduces the number of causes producing detectable effects, and the causes that produce them do it over a smaller TOF range. Figure 8 shows the simulated and reconstructed TOF spectra (top panel) and the resulting neutron energy distribution (bottom panel) after applying a detection threshold of $30$ keVee, corresponding to a neutron energy of $\sim 300$ keV. The value corresponds to the real threshold of MONSTER in the experiment [30]. As can be seen in the figure, the inclusion of the threshold limits the lower neutron energy that can be detected, and hence the original neutron spectrum is never going to be reproduced below that limit. The reconstructed energy distribution has been normalized to the estimated true total number of events $\hat{N}_{true}$, as calculated from Equation 5, and the total number of decays $N_{obs}$.
@@ 4 Conclusions
The methodology discussed in this paper has been applied successfully to obtain the $\beta$-delayed neutron energy spectra emitted in the decays of 85,86As. These measurements served as an experimental validation of the methodology and for producing new physics results. These results will be published in a forthcoming paper.